OpenStudy (anonymous):

If m∠1+m∠2=90 and m∠2=m∠3 then m∠1=m∠3=90 What property is this

OpenStudy (campbell_st):

ummm U think that this is a little wrong it should just finish m<1 + m< 3 = 90

OpenStudy (anonymous):

This is the same question that was asked here: http://openstudy.com/study#/updates/508e0e1ee4b078e4677b70ca ...Where I provided a proof as to why this question does not make sense.

OpenStudy (anonymous):

What does the answer say?

OpenStudy (campbell_st):

well lets look at it this way let m<1 = 70, then m<2 = 20 m<1 + m<2 = 90 true... as written then m<2 = m< 3 so m<2 = 20... so m<3 = 20 and your last statement is m<1 = m<3 = 90........ m<1 = 70 and m<3 = 20... there are not equal and don't equal 90. so there appears a typo...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Okay here I will type it out one more time and will double check it before I post it and tell me if I changed something then

OpenStudy (anonymous):

If M∠1+M∠2=90 and M∠2=M∠3 then M∠1+M∠3=90

OpenStudy (campbell_st):

see how thats different to your posted question.

OpenStudy (anonymous):

I just don't see how a + and a = can really change the property of it but I guess thats why Im asking the question and your the one answering them...Thanks for your time anyway and I will admit yes I did have one little typo and as I said before I apologize for sounding a little rude but I was really frustrated as it is.